Apologies for absence
Alastair and Alison had sent their apologies.
minutes of the last meeting.
These were accepted.
Report from Computing Executive Group
This agenda item was not taken.
Reports from units.
Toby circulated an update on the LDAP problems to computing staff yesterday. He is now on holiday for a week but has left instructions on how to fix any problems that are likely to arise with LDAP. The most recent problems with LDAP appear to be load related and frequently occur on machines that are running as condor nodes. In the short term Toby will be looking at testing out the behaviour of clients machines using a remote LDAP server once he gets back from holiday. In the longer term he hopes to restart the development project to investigate OpenLDAP replication/proxy caching.George reported that Toby was slowly making progress on diagnosing the LDAP problems but that it was difficult to reproduce the problem and increasing the level of debugging could actually alter the behaviour significantly thus making it very difficult to figure out what was causing the problem.
The air-conditioning unit in the Buccleuch Place server room had again started to leak water (from last Friday). The heating engineers suspected that the scavenger pump in the base of the unit may not be working but they needed to get hold of an electrician to confirm this. The setting of the room thermostat had been increased from 16°C to 19°C to reduce the load on the air-conditioning unit. Thanks to Gordon and Neil for coming in at the weekend and emptying the bucket that was catching the condensate leak.
George reported that we had received amended dates for the Appleton Tower work completion and movement of people. Staff on level 4 are moving out today and are, for the most part, moving to level 7 (Dave Robertson's group are moving temporarily to the south computing lab on level 5). Level 8 should be handed over next Monday and people will start to move in on 24th July. The building work on levels 3 and 4 should be completed on 27th August.
There is now a new Dell rack in the Appleton Tower server room. A new Lantronix console server box and a networked advanced power distribution block have been installed in this rack. A 10GBit link has been ordered to connect the switch in this rack back to a switch in the communications rack. Anybody wishing to install a machine in this rack should contact George. The console connectors on the back of the Lantronix have an unusual pin-out and the technicians are making up suitable cables for use between server console ports and the Lantronix box.
Neil asked about the power load distribution within the Appleton Tower server room. George said that there was still work that needed to be done to improve the situation. One possible action would be to plug the UPSs directly into the wall sockets if there were any free within reach of the UPSs. Any other alterations to the power distribution would almost certainly require servers to be shut down first and this would need careful forward planning.
Stephen reported that there was now a test kernel facility for FC6 (via the live/test-kernel.h header file). All the standard kernel modules that we know we are using have been tested against it but units need to let the Managed Platform Unit know about any kernel modules that they are using so that these can also be tested.
Lots of work has been done recently on support for wide screen monitors. The work needs to be done to support them under FC5, whereas such monitors are auto-detected under FC6.
Chris has imported all the packages for Scientific Linux 5 and has created rpm package lists for them. Stephen thought that there would probably be a Scientific Linux 5 project proposal brought to the Development Meeting once Alastair was back.
Auto-detection of the HP L1740 flat panel monitors under FC6 is failing and the resolution being used by the X server is defaulting to 800 x 600 pixel. This is being investigated to see whether any alternative auto-detection software can detect it. If nothing can be found to work then we will need to use a new monitor-specific header file for any desktop machine using one of these monitors.
Research and Teaching.
Tim reported that Iain has changed the GridEngine clusters so that they use a separate LDAP server rather than a local one on each node. Iain will monitor the situation to see whether this improves reliability.
We had intended to just ship Java 1.6 with FC6 (one can also compile Java 1.5 code against Java 1.6), but it has emerged that some users still need the Java 1.5 environment and so it will be necessary to ship this in parallel with Java 1.6.
Iain will be trying to package up firefox 2 to ship with FC6 since it does solve the printing problems people have noticed with the default version of firefox shipped with FC6 (firefox 1.5.10).
Archie has made some progress in getting the Xilinx USB kernel module to work under FC6. It will not be possible to use a single server running Scientific Linux 5 since students need the Xilinx module to talk to devices on their own lab machine. If it isn't possible to get Xilinx working adequately under FC6 then we will need to use a Windows environment (possibly dual boot).
The broken colour printer on level 5 of the Appleton Tower has been replaced by a new HP 4700 colour LaserJet printer; it is still accessible as at13c.
Craig believes that they have a solution to the OpenOffice printing under FC6 (it will only use CUPS). They will provide one or two CUPS servers which will talk to the LPRng servers and advertise the existence of the printers that are available. Alison will be handling the LCFG management of the CUPS servers.
Craig said that they were looking for customers to be early adopters of the test Content Management System that they will shortly be introducing based on zope and plone.
Since the last Operational Meeting the User Support Unit had handled 151 new RT tickets (equivalent to about 15 per working day) and resolved 54% of them. There had been a total of 149 tickets (including both new and existing tickets) resolved over the same period.
Ken reported some figures showing resolution rates against age of recent RT tickets:
|Ticket Age D||Percentage Resolved (or rejected)|
|1 week < D < 2 weeks||69%|
|2 weeks < D < 1 month||85%|
|1 month < D < 3 months||93%|
The move of staff in Appleton Tower from levels 4 to 7 is happening today. The Front Line Support Team started preparing desktop machines for the move yesterday afternoon.
The move of staff home directories from NFS to AFS continues; The total number of accounts (not including temporary ones) which have an AFS home directory is now 153.
Roger reported that the DICE newsletter should be published later today.
There was none.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh